A fair dice is cast what’s the probability that

(a)the result is 4

(b)the result is odd







1
Expert's answer
2021-06-30T16:40:04-0400

(a) Since the dice has 6 faces, the total number of outcomes is n=6.

Of all the outcomes, only m = 1 suits us. So, the wanted probability is "p = \frac{m}{n} = \frac{1}{6}" .

Answer: "p = \frac{1}{6}"

(b) Likewise, the total number of outcomes is n=6.

The number of favorable outcomes is m=3 (1, 3 and 5 points).

So, the wanted probability is "p =\frac{1}{2}" .

Answer: "p =\frac{1}{2}"